Jarod is having a problem with rabbits getting in his vegetable garden, so he decides to fence it in. The length of the garden is 12 feet more than 5 times the width. He needs 132 feet of fencing to do the job. Find the length and width of the garden.
Answer by mananth(16946)

Jarod is having a problem with rabbits getting in his vegetable garden, so he decides to fence it in. The length of the garden is 12 feet more than 5 times the width. He needs 132 feet of fencing to do the job. Find the length and width of the garden.
Assume the width to be x
5 times width = 5x.
12 feet more than 5x ,is 5x+12
2x+2(5x+12)
2x+10x+24=132
12x=108
x=108/12
x=9 which is the width
length is 5x+12
5*9 +12=57 which is the length